OpenStudy (calculusxy):

Geometry: Scale Factors (Enlargement) Haley drew a triangle JKL with the points: J(-3, 0) K(-2, 2) L(0, 0) What will be the new points of triangle MNL using a scale factor of 3 and the origin as the center?

OpenStudy (calculusxy):

@skullpatrol @KamiBug @mathstudent55 @AbdullahM

OpenStudy (calculusxy):

Do I just multiply each of the coordinates by 3?

OpenStudy (abdullahm):

We need to multiply by the scale factor which is 3. J(-3,0) * 3 --> J'(x,y) K(-2,2) * 3 --> K'(a,b) L(0,0) * 3 --> L'(0,0) -3 * 3 = x 0 * 3 = y -2 * 3 = a -2 * 2 = b

OpenStudy (abdullahm):

Basically, yes!

OpenStudy (calculusxy):

What about the center being the origin? Do we need to do anything about that?

OpenStudy (abdullahm):

Nope, 0 * 3 = 0 and so the center remains the center.

OpenStudy (calculusxy):

OK. Thank you.

OpenStudy (xapproachesinfinity):

actually the center changes too

OpenStudy (xapproachesinfinity):

draw the picture to understand it

OpenStudy (calculusxy):

I don't understand.

OpenStudy (abdullahm):

I don't think so.

OpenStudy (xapproachesinfinity):



OpenStudy (xapproachesinfinity):

this is your starting triangle

OpenStudy (xapproachesinfinity):

where is the center of that triangle?

OpenStudy (calculusxy):

By the center do you mean the origin?

OpenStudy (xapproachesinfinity):

no the centroid

OpenStudy (xapproachesinfinity):

the problem is telling the center as well not just scaling the the lengths

OpenStudy (calculusxy):